# Preventing Model Collapse Under Overparametrization: Optimal Mixing Ratios for Interpolation Learning and Ridge Regression

### v1

@@ A.3 Bias
By plugging in $A=C$ and $B={\bm{\Sigma}}$ in (Bach, 2024, Eq (3.9)) and noticing that $(I+m(z){\bm{\Sigma}})$ and $\bm{\Sigma}$} are simultaneously diagonalizable (and therefore commute), we get
$z^{2}\operatorname{Tr}[CQ_{n}(z){\bm{\Sigma}}Q_{n}(z)]$ $\sim\,\operatorname{Tr}[C(I+m(z){\bm{\Sigma}})^{-2}{\bm{\Sigma}}]$
$\quad+\operatorname{Tr}[C(I+m(z){\bm{\Sigma}})^{-2}{\bm{\Sigma}}]\cdot\frac{\operatorname{Tr}[{\bm{\Sigma}}^{2}\left(m(z)^{-1}{\bm{\Sigma}}+I\right)^{-2}]}{n-df_{2}(m(z)^{-1})}$
$\sim\operatorname{Tr}[C(I+m(z){\bm{\Sigma}})^{-2}{\bm{\Sigma}}]\cdot\left(1+\frac{df_{2}(m(z)^{-1})}{n-df_{2}(m(z)^{-1})}\right)$
$\sim\operatorname{Tr}[C(I+m(z){\bm{\Sigma}})^{-2}{\bm{\Sigma}}]\cdot\frac{1}{1-n^{-1}df_{2}(m(z)^{-1})}$
@@
We now show that Equation (3.9) holds. Start by noticing that $nA_{z}=Q_{n}(-z)$. Then, equation (A.7) combined with Lemma A.4 tells us that
$\lim_{t\to\infty}B(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})$ $=\left(\frac{n\lambda}{w}\right)^{2}\operatorname{Tr}[\boldsymbol{\beta}\boldsymbol{\beta}^{\top}A_{\lambda/w}{\bm{\Sigma}}A_{\lambda/w}]$
@@
$\lim_{n\to\infty}\operatorname{Tr}[\boldsymbol{\beta}\boldsymbol{\beta}^{\top}(I+m_{1}{\bm{\Sigma}})^{-2}{\bm{\Sigma}}]$ $=b_{\star}\int\frac{x}{(1+m_{1}x)^{2}}dG$
$\text{ and }\lim_{n\to\infty}n^{-1}df_{2}(m_{1}^{-1})$ $=\gamma\int\frac{m_{1}^{2}x^{2}}{(1+m_{1}x)^{2}}dH$ (A.11)
@@
Plugging $H=G=\delta_{\alpha}$ in equation (3.10) and equation (3.9) yields the asymptotic variance and bias for $\widehat{\boldsymbol{\beta}}_{t,\lambda}$ when ${\bm{\Sigma}}=I$. The proof of log-convexity of risk follows from the proof of Theorem 3.3
@@
$\mathbb{E}(\widehat{\boldsymbol{\beta}}_{t}|\mathbf{X})=(\mathbf{X}^{\top}\mathbf{X})^{\dagger}\mathbf{X}^{\top}\Big(w\mathbb{E}(\mathbf{y}_{t}|\mathbf{X})+(1-w)\mathbb{E}(\widetilde{\mathbf{y}}_{t}|\mathbf{X})\Big)$ (A.13)
@@
$B(\widehat{\boldsymbol{\beta}}_{t};\boldsymbol{\beta})=\boldsymbol{\beta}^{\top}P_{\mathbf{X}}{\bm{\Sigma}}P_{\mathbf{X}}\boldsymbol{\beta}$
where $P_{\mathbf{X}}=I-(\mathbf{X}^{\top}\mathbf{X})^{\dagger}(\mathbf{X}^{\top}\mathbf{X})$. Using (Hastie et al., 2022, Theorem 2), we obtain the analytic value of the bias.
@@ Appendix B Proofs of remaining results
$-\frac{m^{\prime}(z)}{m^{2}(z)}+1$ $=\gamma\int\frac{-x^{2}m^{\prime}(z)}{(1+xm(z))^{2}}dH$ (B.1)
@@
$=\frac{1}{\lambda}\left(f(\lambda/w)-f(\lambda/2-w)\right)$ (B.4)
@@
$\gamma\int\frac{m_{1}^{2}x^{2}}{(1+m_{1}x)^{2}}dH$ $=1-\frac{m_{1}^{2}}{m^{\prime}(-\lambda/w)}$
@@
If ${\bm{\Sigma}}=\alpha I$, we have $G=H$, and we obtain the desired conclusion by the argument above.
@@
Since $\phi_{2}(m_{1})$ is also decreasing and log-convex as function of $w$, this implies that $\lim_{t\to\infty}\lim_{n\to\infty}B(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})$ is decreasing and log-convex in $w$. Therefore, asymptotic generalization error has a unique minima $w^{\star}$. To see the properties of $w^{\star}$, note that by equation (B.9), we have $w^{\star}\geq 1/2$. Using equation (B.8), we obtain that $w^{\star}\rightarrow 1/\varphi$ as $\lambda\rightarrow 0+$. Finally similar to the proof of Theorem 3.2, we have $\mathcal{V}_{\lambda}\rightarrow 0$ as $\lambda\rightarrow\infty$ and $\liminf\limits_{\lambda\rightarrow\infty}\mathcal{B}_{\lambda}>0$. Since $\mathcal{B}_{\lambda}$ is decreasing, we again have $\lim_{\lambda\rightarrow\infty}w^{\star}(\lambda)=1$. This completes the proof of Proposition 3.3.
@@
Recall, the definition $m(z)=\mathbb{E}(A-z)^{-1}$ for some non-negative random variable $A$. Using the definition above, the map $z\to m(-z)$ is a Stieltjes function. Further $\psi(z)=1/z$ is a Stieltjes function by Definition B.1. Using (Schilling et al., 2009, Theorem 6.2(ii) and Corollary 7.9), we obtain that 1/(zm(z))1/(zm(-z)) is a Stieltjes function. Note that 1/(zm(z))=(f(z)/z)11/(zm(-z))=(f(z)/z)-1. Hence, by Definition B.1, we have
@@ Appendix C Dynamic mixing
$w^{\star}_{t}=\operatorname*{arg\,min}_{w}V(\boldsymbol{\beta}_{t};\boldsymbol{\beta}),\quad{\sf Cov}[\boldsymbol{\beta}_{t}|\mathbf{X}]=\sigma^{2}w^{\star}_{t}(\mathbf{X}^{\top}\mathbf{X})^{\dagger}{\bm{\Sigma}}.$ (C.3)
@@
${\sf Cov}[\widehat{\boldsymbol{\beta}}_{t}|\mathbf{X}]$ $=(\mathbf{X}^{\top}\mathbf{X})^{\dagger}\mathbf{X}^{\top}(w^{2}_{1}\sigma^{2}I+2(1-w_{1})^{2}\sigma^{2}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{\dagger}\mathbf{X}^{\top})\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{\dagger}$
@@
Hence the variance is minimized at $w^{\star}_{1}=2/3=(1+w^{\star}_{0})/(2+w^{\star}_{0})$. Further, ${w^{\star}_{1}}^{2}+(1-w^{\star}_{1})^{2}=w^{\star}_{1}$ proving equation (C.3) for $t=1$. Now, for any $t>1$ and any mixing proportion $w_{t}$, we have by induction hypothesis
${\sf Cov}[\widehat{\boldsymbol{\beta}}_{t}|\mathbf{X}]=(\mathbf{X}^{\top}\mathbf{X})^{\dagger}\mathbf{X}^{\top}(w^{2}_{t}\sigma^{2}I+(1-w_{t})^{2}{\sf Cov}(\widetilde{y}_{t-1}|\mathbf{X}))\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{\dagger}$
$=w^{2}_{t}\sigma^{2}(\mathbf{X}^{\top}\mathbf{X})^{\dagger}+(1-w_{t})^{2}(\mathbf{X}^{\top}\mathbf{X})^{\dagger}\mathbf{X}^{\top}\left[\mathbf{X}{\sf Cov}(\widehat{\boldsymbol{\beta}}_{t-1}|\mathbf{X})\mathbf{X}^{\top}+\sigma^{2}I\right]\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{\dagger}$
@@ Appendix D Random Effects
Let ${\bm{\Sigma}}=\sum_{k=1}^{p}s_{k}v_{k}v_{k}^{\top}$ with eigenvalues $s_{1},\dots,s_{p}$ and orthonormal eigenvectors $v_{1},\dots,v_{p}$ Define
$\widehat{H}_{p}(x)=\frac{1}{p}\sum_{k=1}^{p}\mathbf{1}_{s_{k}\leq x},\qquad\widehat{G}_{p}(x)=\frac{1}{\|\boldsymbol{\beta}\|_{2}^{2}}\sum_{k=1}^{p}\langle v_{k},\boldsymbol{\beta}\rangle^{2}\,\mathbf{1}_{s_{k}\leq x}.$
@@
Fix any bounded Lipschitz continuous function $\psi:\mathbb{R}\to\mathbb{R}$ such that $\|\psi\|_{\infty}\leq M$. Set $a_{k}:=\psi(s_{k})$ (which is deterministic given $\bm{\Sigma}$}). Let $\xi_{k}:=\langle v_{k},\boldsymbol{\beta}\rangle$. Since $\beta_{i}$’s are i.i.d. and independent of $\bm{\Sigma}$}, orthonormality gives that, conditional on $\bm{\Sigma}$}, the variables $\xi_{k}$ identically distributed and mutually uncorrelated with $\mathbb{E}\xi_{k}=0$, $\mathbb{E}\xi_{k}^{2}=\tau^{2}$, $\mathbb{E}\xi_{k}^{4}<\infty$. Then
@@
again conditional on $\Sigma$. By characterizations of weak convergence, we have that $\widehat{G}_{p}-\widehat{H}_{p}\Rightarrow 0$ in probability, and thus any weak limit of $\widehat{H}_{p}$ is also a weak limit of $\widehat{G}_{p}$. ∎
@@
$\widehat{G}_{p}(x)=\frac{1}{\|\boldsymbol{\beta}\|_{2}^{2}}\sum_{k=1}^{p}\langle v_{k},\boldsymbol{\beta}\rangle^{2}\,\mathbf{1}_{s_{k}\leq x}=\frac{1}{\|\boldsymbol{\omega}\|_{2}^{2}}\sum_{k=1}^{p}\langle v_{k},\boldsymbol{\omega}\rangle^{2}\,\mathbf{1}_{s_{k}\leq x}=:\widetilde{G}_{p}(x),$

### v2

@@ A.3 Bias
By plugging in $A=C$ and $B={\bm{\Sigma}}$ in (Bach, 2024, Eq (3.9)) and noticing that $(\bm{I}+m(z){\bm{\Sigma}})$ and $\bm{\Sigma}$} are simultaneously diagonalizable (and therefore commute), we get
$z^{2}\operatorname{Tr}[CQ_{n}(z){\bm{\Sigma}}Q_{n}(z)]$ $\sim\,\operatorname{Tr}[C(\bm{I}+m(z){\bm{\Sigma}})^{-2}{\bm{\Sigma}}]$
$\quad+\operatorname{Tr}[C(\bm{I}+m(z){\bm{\Sigma}})^{-2}{\bm{\Sigma}}]\cdot\frac{\operatorname{Tr}[{\bm{\Sigma}}^{2}\left(m(z)^{-1}{\bm{\Sigma}}+\bm{I}\right)^{-2}]}{n-df_{2}(m(z)^{-1})}$
$\sim\operatorname{Tr}[C(\bm{I}+m(z){\bm{\Sigma}})^{-2}{\bm{\Sigma}}]\cdot\left(1+\frac{df_{2}(m(z)^{-1})}{n-df_{2}(m(z)^{-1})}\right)$
$\sim\operatorname{Tr}[C(\bm{I}+m(z){\bm{\Sigma}})^{-2}{\bm{\Sigma}}]\cdot\frac{1}{1-n^{-1}df_{2}(m(z)^{-1})}$
@@
We now show that Equation (3.9) holds. Start by noticing that $n\mathbf{A}_{z}=Q_{n}(-z)$. Then, equation (A.7) combined with Lemma A.4 tells us that
$\lim_{t\to\infty}B(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})$ $=\left(\frac{n\lambda}{w}\right)^{2}\operatorname{Tr}[\boldsymbol{\beta}\boldsymbol{\beta}^{\top}\mathbf{A}_{\lambda/w}{\bm{\Sigma}}\mathbf{A}_{\lambda/w}]$
@@
$\lim_{n\to\infty}\operatorname{Tr}[\boldsymbol{\beta}\boldsymbol{\beta}^{\top}(\bm{I}+m_{1}{\bm{\Sigma}})^{-2}{\bm{\Sigma}}]$ $=b_{\star}\int\frac{x}{(1+m_{1}x)^{2}}dG,$
$\text{ and }\lim_{n\to\infty}n^{-1}df_{2}(m_{1}^{-1})$ $=\gamma\int\frac{m_{1}^{2}x^{2}}{(1+m_{1}x)^{2}}dH.$ (A.11)
@@
Plugging $H=G=\delta_{\alpha}$ in equation (3.10) and equation (3.9) yields the asymptotic variance and bias for $\widehat{\boldsymbol{\beta}}_{t,\lambda}$ when ${\bm{\Sigma}}=\bm{I}$. The proof of log-convexity of risk follows from the proof of Theorem 3.3
@@
$\mathbb{E}(\widehat{\boldsymbol{\beta}}_{t}|\mathbf{X})=(\mathbf{X}^{\top}\mathbf{X})^{\dagger}\mathbf{X}^{\top}\Big(w\mathbb{E}(\mathbf{y}_{t}|\mathbf{X})+(1-w)\mathbb{E}(\widetilde{\mathbf{y}}_{t}|\mathbf{X})\Big).$ (A.13)
@@
$B(\widehat{\boldsymbol{\beta}}_{t};\boldsymbol{\beta})=\boldsymbol{\beta}^{\top}P_{\mathbf{X}}{\bm{\Sigma}}P_{\mathbf{X}}\boldsymbol{\beta},$
where $P_{\mathbf{X}}=\bm{I}-(\mathbf{X}^{\top}\mathbf{X})^{\dagger}(\mathbf{X}^{\top}\mathbf{X})$. Using (Hastie et al., 2022, Theorem 2), we obtain the analytic value of the bias.
@@ Appendix B Proofs of remaining results
$-\frac{m^{\prime}(z)}{m^{2}(z)}+1$ $=\gamma\int\frac{-x^{2}m^{\prime}(z)}{(1+xm(z))^{2}}dH.$ (B.1)
@@
$=\frac{1}{\lambda}\left(f\left(\frac{\lambda}{w}\right)-f\left(\frac{\lambda}{2-w}\right)\right)$ (B.4)
@@
$\gamma\int\frac{m_{1}^{2}x^{2}}{(1+m_{1}x)^{2}}dH$ $=1-\frac{m_{1}^{2}}{m^{\prime}(-\lambda/w)}.$
@@
If ${\bm{\Sigma}}=\alpha\bm{I}$, we have $G=H$, and we obtain the desired conclusion by the argument above.
@@
Since $\phi_{2}(m_{1})$ is also decreasing and log-convex as function of $w$, this implies that $\lim_{n\to\infty}\lim_{t\to\infty}B(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})$ is decreasing and log-convex in $w$. Therefore, asymptotic generalization error has a unique minima $w^{\star}$. To see the properties of $w^{\star}$, note that by equation (B.9), we have $w^{\star}\geq 1/2$. Using equation (B.8), we obtain that $w^{\star}\rightarrow 1/\varphi$ as $\lambda\rightarrow 0+$. Finally similar to the proof of Theorem 3.2, we have $\mathcal{V}_{\lambda}\rightarrow 0$ as $\lambda\rightarrow\infty$ and $\liminf\limits_{\lambda\rightarrow\infty}\mathcal{B}_{\lambda}>0$. Since $\mathcal{B}_{\lambda}$ is decreasing, we again have $\lim_{\lambda\rightarrow\infty}w^{\star}(\lambda)=1$. This completes the proof of Proposition 3.3.
@@
Recall, the definition $m(z)=\mathbb{E}(A-z)^{-1}$ for some non-negative random variable $A$. Using the definition above, the map $z\to m(-z)$ is a Stieltjes function. Further $\psi(z)=1/z$ is a Stieltjes function by Definition B.1. Using (Schilling et al., 2009, Theorem 6.2(ii) and Corollary 7.9), we obtain that 1/(zm(z))1/(zm(-z)) is a Stieltjes function. Note that $(zm(-z))^{-1}=z^{-1}f(z)+1$. Hence, by Definition B.1, we have
@@ Appendix C Dynamic mixing
$w^{\star}_{t}=\operatorname*{arg\,min}_{w}V(\boldsymbol{\beta}_{t};\boldsymbol{\beta}),\quad{\sf Cov}[\boldsymbol{\beta}_{t}|\mathbf{X}]=\sigma^{2}w^{\star}_{t}(\mathbf{X}^{\top}\mathbf{X})^{\dagger}.$ (C.3)
@@
${\sf Cov}[\widehat{\boldsymbol{\beta}}_{t}|\mathbf{X}]$ $=(\mathbf{X}^{\top}\mathbf{X})^{\dagger}\mathbf{X}^{\top}(w^{2}_{1}\sigma^{2}\bm{I}+2(1-w_{1})^{2}\sigma^{2}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{\dagger}\mathbf{X}^{\top})\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{\dagger}$
@@
Hence the variance is minimized at $w^{\star}_{1}=2/3=(1+w^{\star}_{0})/(2+w^{\star}_{0})$. Further, ${w^{\star}_{1}}^{2}+2(1-w^{\star}_{1})^{2}=w^{\star}_{1}$ proving equation (C.3) for $t=1$. Now, for any $t>1$ and any mixing proportion $w_{t}$, we have by induction hypothesis
${\sf Cov}[\widehat{\boldsymbol{\beta}}_{t}|\mathbf{X}]=(\mathbf{X}^{\top}\mathbf{X})^{\dagger}\mathbf{X}^{\top}(w^{2}_{t}\sigma^{2}\bm{I}+(1-w_{t})^{2}{\sf Cov}(\widetilde{y}_{t-1}|\mathbf{X}))\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{\dagger}$
$=w^{2}_{t}\sigma^{2}(\mathbf{X}^{\top}\mathbf{X})^{\dagger}+(1-w_{t})^{2}(\mathbf{X}^{\top}\mathbf{X})^{\dagger}\mathbf{X}^{\top}\left[\mathbf{X}{\sf Cov}(\widehat{\boldsymbol{\beta}}_{t-1}|\mathbf{X})\mathbf{X}^{\top}+\sigma^{2}\bm{I}\right]\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{\dagger}$
@@ Appendix D Random Effects
Let ${\bm{\Sigma}}=\sum_{k=1}^{p}s_{k}\bm{v}_{k}\bm{v}_{k}^{\top}$ with eigenvalues $s_{1},\dots,s_{p}$ and orthonormal eigenvectors $\bm{v}_{1},\dots,\bm{v}_{p}$ Define
$\widehat{H}_{p}(x)=\frac{1}{p}\sum_{k=1}^{p}\mathbf{1}_{s_{k}\leq x},\qquad\widehat{G}_{p}(x)=\frac{1}{\|\boldsymbol{\beta}\|_{2}^{2}}\sum_{k=1}^{p}\langle\bm{v}_{k},\boldsymbol{\beta}\rangle^{2}\,\mathbf{1}_{s_{k}\leq x}.$
@@
Fix any bounded Lipschitz continuous function $\psi:\mathbb{R}\to\mathbb{R}$ such that $\|\psi\|_{\infty}\leq M$. Set $a_{k}:=\psi(s_{k})$ (which is deterministic given $\bm{\Sigma}$}). Let $\xi_{k}:=\langle\bm{v}_{k},\boldsymbol{\beta}\rangle$. Since $\beta_{i}$’s are i.i.d. and independent of $\bm{\Sigma}$}, orthonormality gives that, conditional on $\bm{\Sigma}$}, the variables $\xi_{k}$ identically distributed and mutually uncorrelated with $\mathbb{E}\xi_{k}=0$, $\mathbb{E}\xi_{k}^{2}=\tau^{2}$, $\mathbb{E}\xi_{k}^{4}<\infty$. Then
@@
again conditional on $\bm{\Sigma}$}. By characterizations of weak convergence, we have that $\widehat{G}_{p}-\widehat{H}_{p}\Rightarrow 0$ in probability, and thus any weak limit of $\widehat{H}_{p}$ is also a weak limit of $\widehat{G}_{p}$. ∎
@@
$\widehat{G}_{p}(x)=\frac{1}{\|\boldsymbol{\beta}\|_{2}^{2}}\sum_{k=1}^{p}\langle\bm{v}_{k},\boldsymbol{\beta}\rangle^{2}\,\mathbf{1}_{s_{k}\leq x}=\frac{1}{\|\boldsymbol{\omega}\|_{2}^{2}}\sum_{k=1}^{p}\langle\bm{v}_{k},\boldsymbol{\omega}\rangle^{2}\,\mathbf{1}_{s_{k}\leq x}=:\widetilde{G}_{p}(x),$